OpenStudy (anonymous):

find the diameter of a circle whose circumference is 62.8 feet possible answers: 18,22,20,21

OpenStudy (elonasushchik):

You have the answer. I don't want to get attacked again! Lol

OpenStudy (anonymous):

oh lol by possible I ment abcd type possible

OpenStudy (anonymous):

r^2 * pi (3.14) = the circumference, correct? 62.8 / 3.14 = r^2 r is exactly half of a circle. so * it by two to get the diameter <33 or, you can plug in the answers to this: ? / 2 = ?? ??^2 * 3.14

OpenStudy (anonymous):

well 33x2 so 66 ?

OpenStudy (elonasushchik):

Correct !

OpenStudy (mrnood):

that is NOT correct

OpenStudy (anonymous):

but that's not an answer choice

OpenStudy (anonymous):

not correct. that is not an answer choice.

OpenStudy (mrnood):

@silverr the formula you gave for the circumference is also incorrect

OpenStudy (elonasushchik):

@MrNood hello?!

OpenStudy (anonymous):

i kinda thought so. i haven't done this for a while. what is the formula ?

OpenStudy (anonymous):

I don't remember sorry

OpenStudy (mrnood):

Oh come ON - you are studying circles: Area = ? (it has pi and r^2 in it) Circumference = ? (it has pi and r in it)

OpenStudy (catlover5925):

@ElonaSushchik has been reported several times for giving wrong answers, plz stop telling people they are correct with out doing the work first @ElonaSushchik @Abhisar

OpenStudy (anonymous):

seriously?

OpenStudy (catlover5925):

yes i am not kidding

OpenStudy (anonymous):

oh,, i see. that is no good, shame.

OpenStudy (mrnood):

SO @killer2490 Write the formula for circumference (it has pi and r in it)

OpenStudy (anonymous):

ahh well please im really pressed for time and have a lot of work to do I know rules are rules but could you guys just help me out and skip explanations please

OpenStudy (anonymous):

OH WAIT i see,,, i was giving the area formula! circumference is 2*pi*r, correct?

OpenStudy (mrnood):

NO This is trival question and an important tjhing for you to know just write the formula

OpenStudy (anonymous):

so demanding.... you mind moving on to a different person not comfortable with your help

OpenStudy (mrnood):

It is against the rules to ask for the answer. I will help if oyu are willing to make any attempt for yourself

OpenStudy (anonymous):

if I were capable of obtaining my own answer I wouldn't be here asking

OpenStudy (mrnood):

you are capable you just need the right guidance. It is against the rules to ask for th eanswer as I said before

OpenStudy (anonymous):

no arguments, please,, it's okay ^^

OpenStudy (anonymous):

look man I was special ed ever since I was a kid I have a 20 sec memory span and can don't much for my self they never put me in a math until 9th grade thinking I couldn't handle it and they were right so if your not gonna spare me the details alright I guess

OpenStudy (mrnood):

so if I tell you the answer is 18 - how do you know that I know any better than you?

OpenStudy (anonymous):

abandoned likr always X/

OpenStudy (anonymous):

Circumference is 2 * pi * r Circumference is pi*diameter C = pi*D D = C/pi D = 62.8/3.14 D = 20

OpenStudy (mrnood):

Rubbish- not abandoned - I am offering help BUt spoon feeding you the answer will leave you dependant for the rest of your life man up - try a little effort - it works

OpenStudy (anonymous):

thanks silver

OpenStudy (anonymous):

you understand, though, right ?

OpenStudy (anonymous):

eah no but its cool not going to colledge im a military man XD so forget geo

OpenStudy (anonymous):

can i explain it to you?

OpenStudy (anonymous):

its alright man remember 20 sec span so it wouldn't stick anyway

OpenStudy (catlover5925):

i think @silverr is a woman??

OpenStudy (mrnood):

I fear for th eworld that they would put a gun in your hand

OpenStudy (anonymous):

ah,, okay.. are you sure ? you will need it for school...

OpenStudy (anonymous):

im 12th grade schools pretty much done

OpenStudy (anonymous):

nood....quite the spiteful one arnt you

OpenStudy (anonymous):

yes, yes, i am a woman, thank you for that. they said otherwise ? not done yet,,, though.

OpenStudy (mrnood):

Not spiteful - just despairing of those who will not work for themself

OpenStudy (anonymous):

ahh well ill leave you with that comment and silver a girl who likes anime high five! XD

OpenStudy (catlover5925):

@abram134

OpenStudy (anonymous):

@mrnood so would you rather have him filing your taxes???

OpenStudy (anonymous):

@duckyblossom why are you even commenting on this bro